Is there any lower limsit to the electric force between twon particles placed at a sepration of 1 cm?

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Electric force between two particles separated at a definite distance is given by Coloumb’s law and so it is definite. However, in the sense of choosing the charges themselves, since there is lower limit on the charge that exists(quantization of charge) equal toe=1.6\times 10^{-19}C, the smallest force that can exist is when charges with chargeeand is equal to\dfrac{1}{4\pi\epsilon_0}\dfrac{e^2}{R^2}.